solving for zeros the equation is. y=x^(4)-13x^(2)+36
Answer by stanbon(75887)

solving for zeros the equation is. y=x^(4)-13x^(2)+36
-------------------
x^4 - 9x^2-4x^2 + 36 = 0
---
x^2(x^2-9) -4(x^2-9) = 0
---
(x^2-9)(x^2-4) = 0
---
(x-3)(x+3)(x-2)(x+2) = 0
=========
Zeroes: 3,-3,2,-2
=======================
Cheers,
Stan H.